Last year, Big W Company reported earnings per share of $2.20 when its stock was selling for $55.00. If its earnings this year increase by 10 percent and the P/E ratio remains constant, what will be the price of its stock? (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
Answer #1

Earnings per share of last year = $2.20

Market price per share in the last year = $55

P/E ratio of last year = Market price per share in the last year/Earnings per share of last year

= 55/2.20

= 25

Increase in earnings per share this year = 10%

= 2.20 x 10%

= $0.22

Expected Earnings per share this year = Earnings per share of last year + Increase in earnings per share this year

= 2.20 + 0.22

= $2.42

P/E ratio remains constant in the current year.

Price of the stock in the current year =  P/E ratio x Expected Earnings per share this year

= 25 x 2.42

= $60.5

Please ask if you have any query related to the question. Thank you